Two mirrors are placed at a 90° angle to each other. A light ray strikes one mirror 0.620 m from the intersection of the mirrors with an incident angle of 36.5°. The ray then travels a distance d before reflecting from the second mirror. (a) What is the distance d? answer should be in meters.
1
Expert's answer
2011-06-20T13:03:27-0400

Let A be point of intersection of 2 mirrors. A light ray strikes one mirror at point O.
OA = a = 0.62 m,& α = 36.5. So
cosα = a/d
d = a/cosα = 0.62/cos36.5 =& 0.62/0.803 =& 0.77 m.